OpenStudy (anonymous):

Of the 250 people at a concert, 160 are men. Half of the rest of the people are children. What percent of the people are women? A. 9% B. 16% C. 18% D. 36%

OpenStudy (anonymous):

250 - 160 = 90 90/2 = 45 45/250 = .18 .18*100 = 18%